OpenStudy (spring98):

Which shows the expressions in the order they would appear on a number line from least to greatest?

OpenStudy (spring98):

which of theses link is from least to greatest

OpenStudy (spring98):

@satellite73 @jim_thompson5910 @nincompoop @nonon @horsegirl27 @Asad786

OpenStudy (spring98):

plz help

jimthompson5910 (jim_thompson5910):

use a calculator to convert each number to decimal form for example 4^3 = 4*4*4 = 16*4 = 64

OpenStudy (spring98):

6.3 and 3.6 and1.5

OpenStudy (spring98):

that is the calculated solution with urs on top of it

jimthompson5910 (jim_thompson5910):

how about the square root of 7 ?

OpenStudy (spring98):

oh so would the answer choice be B?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (spring98):

is it right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (spring98):

thanks